# Artificial Intelligence Exploring the Patent Field

### v1

@@ Abstract
Advanced language-processing and machine-learning techniques promise massive efficiency improvements in the previously widely manual field of patent and technical knowledge management. This field presents large-scale and complex data with very precise contents and language representation of those contents. Particularly, patent texts can differ from mundane texts in various aspects, which entails significant opportunities and challenges. This paper presents a systematic overview of patent-related tasks and popular methodologies with a special focus on evolving and promising techniques. Language processing and particularly large language models as well as the recent boost of general generative methods promise to become game changers in the patent field. The patent literature and the fact-based argumentative procedures around patents appear almost as an ideal use case. However, patents entail a number of difficulties with which existing models struggle.
The paper introduces fundamental aspects of patents and patent-related data that affect technology that wants to explore or manage them. It further reviews existing methods and approaches and points out how important reliable and unbiased evaluation metrics become. Although research has made substantial progress on certain tasks, the performance across many others remains suboptimal, sometimes because of either the special nature of patents and their language or inconsistencies between legal terms and the everyday meaning of terms. Moreover, yet few methods have demonstrated the ability to produce satisfactory text for specific sections of patents. By pointing out key developments, opportunities, and gaps, we aim to encourage further research and accelerate the advancement of this field.
@@ 1 Introduction
The recent presentation of powerful general foundation models, and previously unthinkably large models for generative tasks and particularly language processing have made widely sensed waves also in the general public. These tools appear to be a practically ideal match for one of the main resources of technology in our society, namely the patent literature. Despite rapid developments, patents nowadays still mainly rely on text and drawings to capture the essence of an invention and disclose them to the public. Whereas already individual patents describing small details that often miss the big picture are indigestible for most due to the exclusive use of an accurate text description and figures, the entire patent literature due to its size definitely is also for experts.
@@
Letters patents or short patents as a form of intellectual property grant the holder temporary rights to suppress competing use of an invention in exchange for a disclosure of the invention. It was once established to promote and/or control technical innovation and progress [1]. The surge in global patent applications as well as the rapid technological progress forming the state of the art pose a formidable challenge to patent offices and related practitioners. Traditional manual methods of patent drafting and analysis are increasingly inadequate for managing, processing, and analyzing large-scale and complicated patent data. Therefore, advanced computational techniques for the automation of patent tasks are attractive to improve efficacy of both the patent and intellectual property management process itself in corporations and patent offices but also for extracting information from this vast knowledge base [2].
Artificial intelligence (AI) has been evolving at an unprecedented pace, leading to the emergence of various models and methodologies that command skills previously considered available only to the human mind. Notably, transformer-based large models have dramatically affected many fields, such as natural language processing (NLP) [3] and computer vision (CV) [4]. This sudden growth in performance promises significant advancements in the patent domain over the previously possible [5]. The diversity of the tasks and needs of the patent field and the many new technologies and tools on the other side cries for a review of the latest developments in artificial intelligence in the patent field and to uncover the potential of latest techniques.
@@
Previous surveys reported the early steps of smart and automated methods for patent analysis [2], first deep learning methods opening a wider range of still simpler patent tasks [5], or one specific patent task, such as patent retrieval [6]. The recent milestones in machine learning, such as transformers and achievements in language models could not have been predicted. Accordingly, we aim to delineate an overview of all patent-related tasks and popular methodologies with a special focus on most recent or even evolving techniques. We want to include the two applications of patent analysis and patent generation tasks (Figure 1). While analysis tasks focus on understanding and usage of patents, generation tasks involve automatically generating patent texts.
Figure 1: Patent-related tasks.
We want to present a highly auspicious but widely neglected field to the community. Therefore, we provide a systematic overview of automated patent tasks, including fundamental concepts, datasets, methodologies, tasks, and future research directions to serve as a reference for both novices and experts. Specifically, we list the main contributions as follows.
  1. (1)
    We provide an in-depth introduction to the foundational aspects of patents in Section 3, including the composition of patent documents and patent life-cycle. Additionally, we describe related surveys to contextualize the paper. We intend to demonstrate the background knowledge in the field for readers unfamiliar with patents.
  2. (2)
@@
  3. (3)
    We conduct a technique-oriented review in the patent domain in Section 5, aiming to provide a clear roadmap from the methodology perspective. We demonstrate various patent representations, including text, numerical, graph, and image features. In addition, we introduce different types of models, ranging from formula-based to deep learning-based approaches. Moreover, we summarize some useful methods and algorithms that can be integrated to improve performance, such as multimodal models and domain adaptation.
  4. (4)
    We list common evaluation metrics for patent tasks in Section 6, including metrics for classification, information retrieval, and text generation. Readers can compare the performance of different approaches or analyze the strengths and limitations of various methods by understanding the evaluation metrics.
  5. (5)
    We summarize nine patent analysis tasks (subject classification, patent retrieval, information extraction, novelty prediction, granting prediction, litigation prediction, valuation, technology forecasting, and innovation recommendation) in Section 7 and four generation tasks (summarization, translation, simplification, and patent writing) in Section 8. Specifically, we systematically demonstrate task definitions and relevant methodologies in detail. We show a comprehensive yet accessible insight into this area, enabling readers to grasp the nuances of the field more effectively.
  6. (6)
@@
2 What is already possible and what is not?
At first sight, recent large models for generative tasks and language processing appear to be a perfect match for the patent literature and the highly text-based processes flanking patents. Despite this seemingly perfect match for knowledge extraction and language processing, the field of machine learning on patents and patent-related aspects is still underdeveloped and not highly prominent yet. Previous research includes analysis and data extraction as well as automation of procedures. However, missing benchmark tests, e.g., a reference data set, and established metrics complicate performance rating and comparisons if several different methods aim to solve the same tasks. Furthermore, a major advantage of the patent literature and related documents is their—on average—high language precision and accuracy as well as the availability. Still, many studies use closed-source data for training and evaluation.
Patent documents comprise multiple parts that can be used for patent analysis, including texts, drawings, and additional information, such as subject classification codes and the office correspondence for establishing a patent, e.g., the examination, the maintenance, and potential litigation. Accordingly, researchers leverage different approaches to process different parts for analysis. Text is typically the dominant part of patents. Although transformer-based large language models have demonstrated enormous potential in the general domain, the patent field seems to be behind and most studies investigate older models, such as simple parsers or embedding with subsequent more conventional statistical and classification methods. These models have once demonstrated good performance but have recently been outperformed by transformer-based models across various text-based tasks. The largest models based on transformers adapted and tested in the patent domain are on the level of GPT-2 [7] with approximately 1.5 billion parameters and therefore by a factor of at least 100 smaller than more recent models, such as switch transformers [8] that scale the model to trillion parameters.
Moreover, most research has focused on short texts, such as abstracts and claims, while ignoring long description, which contains detailed and specific information about an invention. The reason is simple: models that can process the text of an entire description or even several patents had not been available or hardly manageable. If the text is tokenized as common for language by now, the average number of tokens per patent description of a recently compiled patent dataset exceeds 11,000 [9]. In contrast, the input context limit of most language models is 2,048, while larger or specifically designed models, such as Longformer [10], can increase the input length to 4,096. Notably, the recent GPT-4 model has supported the context length to 32,768 tokens [^1] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo, which appears promising to process long patent descriptions. However, titles and abstracts of patents can be surprisingly generic. Therefore, using descriptions that provide comprehensive details and specific embodiment of the invention for patent analysis is highly important, which is neglected by most of the current research.
Text generation tasks are less investigated due to the level of complexity of patents and patent-related texts. The information density and need for precision and accuracy is higher than in everyday language and particularly of the majority of tasks currently tested with large language models. Instead of the empty verbiage that many of the latter still generate, patent texts however are highly technical and must be precise. However, due to the need of large quantities of data, the majority of the training corpora for most of the pre-trained language models might have to be colloquial and relatively sloppy with a minority from belletristic texts, which may have higher quality but often swap precision and accuracy for style and language originality.
Certain very specific tasks, such as automated translation, have been deeply studied and brought to a high level of quality so that they have been in production at offices, e.g., even available to casual users at the European Patent Office. The obvious and important patent writing tasks, however, are far from satisfactory, such as patent application drafting, e.g., the description or the claims, the examination, and formulating correspondence. Although some studies have proposed methods to generate different sections of patents, there is no acceptable solution or result from even a precursor for any of the tasks so far.
Apart from patent texts, researchers also use metadata and images for patent analysis and particularly linking files and fields. Metadata typically include citations/references, classification codes, inventors, applicants, assignees, law firms, and examiners. Researchers have exploited the relationship information among patents through citations and classification codes based on graph neural networks for patent classification and patent valuation. Solely relying on classification codes can sometimes be misleading and become a problem because patents may span multiple technology areas. Furthermore, the publicly used classification is rather coarse and notably coarser than what offices use internally for their search or assigning competent examiners. Importantly, metadata provide high-level information but do typically not include details of the invention, which limits the possibility of capturing the invention and certainly not technical nuances.
The images, called drawings or figures in patents, could leverage the development of image processing techniques, particularly convolutional neural networks. However, the figures can be very generic, particularly in method inventions. The fact that figures are not supposed to contain (significant) text, can have technological advantages. Patent figures from computer aided design (CAD) models may be harder to match with corresponding ones from other documents due to the high number of lines representing details, which tend to overwhelm convolutional neural networks. However, many patents and patent applications contain simpler sketches and schematic figures with the intention to not disclose more than necessary. Such simplifications can be beneficial for processing with artificial intelligence. By now, researchers have applied convolutional neural networks in patent subject classification and patent retrieval. However, considering the above, the performance of patent classification based on images only appears to be much worse than using texts. Therefore, hybrid methods that leverage multiple sources for patent analysis, such as texts, images, and classification codes, appear promising.
@@ 3 Brief Background
In this section, we want to give those readers that are not familiar with the patent systems a quick introduction to the background. Firstly, we illustrate the composition of patent documents in Section 3.1. Secondly, we summarise the patent life cycle in Section 3.2, including pre-grant and post-grant stages.
@@ 3.1 Patent Document
Patent documents are the central pieces for the protection of intellectual property and document inventions. Patents require applicants and/or inventors to publicly disclose their inventions while securing exclusive rights to obtain benefits. Patent documents describe in detail new inventions and delineate the scope of patent rights granted to patent holders. These documents are key parts of the patenting process and are publicly accessible once patents are issued. The format and content can vary by jurisdiction but normally include the following elements. An example patent document is shown in Figure 2.
Figure 2: An example of granted patent document (US Patent 11,824,732 B2).
Publication information includes patent number and date of patent publication.
@@
Citations are lists of prior arts and other patents referenced in the document.
@@
Drawings are visual representations of the invention, disclosing important aspects of the invention as well as embodiments to support the textual description.
@@
Claims define the legal boundaries of the patent. Each claim is a single sentence describing a specific aspect of the invention and its uniqueness.
@@ 3.2 Patent Life Cycle
The patent life cycle encompasses several stages, from the initial conception of an invention to its eventual expiration. It can be broadly divided into pre-grant and post-grant phases, as shown in Figure 3.
Figure 3: Patent life-cycle from pre-grant to post-grant stage.
Pre-grant Stage. In the beginning, inventors conceptualize, design, and develop inventions. If inventors hope to obtain patent protection for their inventions, they need to make applications to the patent office. To ensure novelty, inventors thoroughly search for existing patents and public disclosures. Additionally, patent documents need drafting, describing the invention in detail, including its specifications, claims, abstract, and accompanying drawings. The drafting process typically requires the expertise of a patent professional, such as a patent engineer, attorney, or agent. After the patent application is submitted to the patent office, examiners at the office will screen and evaluate the application for compliance with formal, legal, and technical requirements. This examination typically involves correspondence of the examiner with the inventor or their representation, where clarifications, amendments, or arguments are submitted. Notably, an assessment will determine if the invention meets the criteria for patentability, including novelty, inventiveness (non-obviousness), and typically less strictly commercial utility. Upon examination and potential resolution of any objections, the patent is either granted rarely in its original, more frequently in a restricted form or rejected. If the original priority date should be maintained, the invention must not be extended during the examination.
Post-grant Stage. The granted patent is published, disclosing the details of the invention to the public. Classification codes are given to granted patents for easy search and management. Maintenance fees should be paid regularly to keep the patent in force. The patent owner can enforce patent rights through legal action when infringement occurs. In addition, third parties can question the validity of patents, and patent owners need to defend the challenges. Besides, companies can analyze patents for technology insights and make strategic plans. Finally, the invention enters the public domain after the expiration of patents, allowing anyone to use it without infringement.
@@ 3.3 Related Surveys
Patent Analysis. Various surveys in the past summarized certain aspects of patents from a knowledge or procedural perspective and patent analysis. The work of Abbas et al. represents early research on patent analysis [2]. These early methods included text mining and visualization approaches, which paved the way for future research. Deep learning for knowledge and patent management started more recently [5]. Krestel et al. identified eight types of sub-tasks attractive for deep learning methods, specifically supporting tasks, patent classification, patent retrieval, patent valuation, technology forecasting, patent text generation, litigation analysis, and computer vision tasks. Some surveys reviewed specific topics of patent analysis. Patent retrieval, for example, i.e., the search for relevant patent documents, may appear similar to web search but has substantially different objectives and constraints, to which language processing and machine learning need to adapt [6].
@@
Patent Usage. The patent literature documents major technological progress and constitutes a knowledge database. Furthermore, it is known that the translation of methods of one domain to another or the combination of techniques or technology from various fields can lead to major innovation itself. As such, the contents of the patent literature appear highly attractive for systematic analysis for identifying patterns but also generatively as a tool supporting design. However, the language of modern patents has substantially evolved and diverged from normal technical writing so that recent patent documents are typically hardly digestible for the normal reader and also contain deceptive elements added for increasing the scope or camouflage important aspects. Appropriate data science techniques need to consider such aspects to mine patent databases for engineering design [11]. If they do, data science applied to this body of codified technical knowledge can inform design theory and methodology but also form design tools and strategies. Particularly in early stages of innovation, language processing can help with ideation, trend forecasting, or problem–solution matching [12].
@@
Patent-Related Language Generation. Whereas the above-listed techniques harvest the existing patent literature for external needs, the strong text-dependence of the patent field suggests the use of language models in a generative way. Language processing techniques can, for instance, translate the peculiar patent language to more understandable technical texts, summarize texts, or generate patent texts based on prompts [13]. The recent rapid development of generative language models, especially large language models (LLMs) [14] may stimulate more patent-related generation tasks.
@@
Related Topics. A variety of further tasks and applications still have close ties to the patent domain or encompass it together with other fields. Intellectual property (IP), for instance, includes copyrights, trade marks, designs, and a number of more special rights beyond patents. The wider scope beyond technical inventions obviously affects the tools. The combined intellectual property literature allows concordant knowledge management, technology management, economic value estimation, and information extraction as well as management [15]. The patent field has further close ties to the general legal domain and shares procedural aspects as well as the precision of language. Likewise, most patent professionals have substantial legal training or a legal degree as also a large share of their work involves the legal side. Accordingly, natural language processing techniques in the more general field of law may influence patent-related techniques in the future [16].

### v2

# Natural Language Processing in Patents: A Survey

@@ Abstract
Patents, encapsulating crucial technical and legal information, present a rich domain for natural language processing (NLP) applications. As NLP technologies evolve, large language models (LLMs) have demonstrated outstanding capabilities in general text processing and generation tasks. However, the application of LLMs in the patent domain remains under-explored and under-developed due to the complexity of patent processing. Understanding the unique characteristics of patent documents and related research in the patent domain becomes essential for researchers to apply these tools effectively. Therefore, this paper aims to equip NLP researchers with the essential knowledge to navigate this complex domain efficiently. We introduce the relevant fundamental aspects of patents to provide solid background information, particularly for readers unfamiliar with the patent system. In addition, we systematically break down the structural and linguistic characteristics unique to patents and map out how NLP can be leveraged for patent analysis and generation. Moreover, we demonstrate the spectrum of text-based patent-related tasks, including nine patent analysis and four patent generation tasks.
Natural Language Processing in Patents: A Survey
Lekang Jiang  and Stephan Goetz University of Cambridge {lj408, smg84}@cam.ac.uk
@@ 1 Introduction
Patents, a form of intellectual property (IP), grant the holder temporary rights to suppress competing use of an invention in exchange for a complete disclosure of the invention. It was once established to promote and/or control technical innovation and progress (Frumkin, 1947). The surge in global patent applications and the rapid technological progress pose formidable challenges to patent offices and related practitioners (Krestel et al., 2021). These challenges overwhelm traditional manual methods of patent drafting and analysis. Consequently, there is a significant need for advanced computational techniques to automate patent-related tasks. Such automation not only enhances the efficiency of patent and IP management but also facilitates the extraction of valuable information from this extensive knowledge base (Abbas et al., 2014).
Researchers have investigated machine learning (ML) and natural language processing (NLP) methods for the patent field with highly technical and legal texts (Krestel et al., 2021). In addition, the recent large language models (LLMs) have demonstrated outstanding capabilities across a wide range of general domain tasks (Zhao et al., 2023; Min et al., 2023). These models are promising to become valuable tools in managing and drafting patent literature, the crucial resource that documents technological advances.
However, compared to the significant success of LLMs in the general domain, the application of LLMs in patent-related tasks remains under-explored due to the texts’ and the field’s complexity. NLP researchers need to deeply understand the unique characteristics of patent documents to develop useful models for the patent field. Therefore, we aim to equip NLP researchers with the essential knowledge by presenting this highly auspicious but widely neglected field to the NLP community.
Figure 1: Patent-related tasks
@@
Previous surveys reported the early stages of smart and automated methods for patent analysis (Abbas et al., 2014), first deep learning methods, which opened a wider range of still simpler patent tasks (Krestel et al., 2021), or specific individual patent tasks, such as patent retrieval (Shalaby and Zadrozny, 2019). The recent advancements in NLP were unforeseen, particularly the performance boost when language models are scaled up. Accordingly, we specifically delineate a survey of popular NLP methodologies for patent documents with a special focus on the most recent and evolving techniques. We have included the two applications of patent analysis and generative patent tasks (Fig. 1). Whereas analysis focuses on understanding and usage of individual patent documents or a group of patents, generation tasks aim at automatically generating patent texts.
We provide a systematic survey of NLP applications in the patent domain, including fundamental concepts, insights on patent texts, development trends, datasets, tasks, and future research directions to serve as a reference for both novices and experts. Specifically, we cover the following topics:
  1. 1.
    We provide an introduction to the fundamental aspects of patents in Section 2, including the composition of patents and the patent life-cycle. This section is particularly for those readers still unfamiliar with it and a refresher for others.
  2. 2.
    We analyze the unique structural and linguistic characteristics of patent texts with NLP techniques to process these texts in Section 3. Readers will readily understand the challenges of automated patent processing and the development trends of NLP in the patent field.
  3. 3.
@@
  4. 4.
    We evaluate nine patent analysis tasks (subject classification, patent retrieval, information extraction, novelty prediction, granting prediction, litigation prediction, valuation, technology forecasting, and innovation recommendation) in Section 5 and four generation tasks (summarization, translation, simplification, and patent writing) in Section 6. Specifically, we systematically demonstrate task definitions and relevant methodologies in detail. We show a comprehensive yet accessible insight into this area, enabling readers to grasp the nuances of the field more effectively.
  5. 5.
@@ 2 Brief Background
Figure 2: An example of granted patent document (US Patent 11,824,732 B2). This is a US patent, and other countries have their own patent systems with largely similar requirements.
@@ 2.1 Patent Document
Patent documents are central elements for the protection of intellectual property and further document inventions. Patents require applicants and/or inventors to publicly disclose their inventions in detail to secure exclusive rights and obtain benefits in return. Patent documents describe new inventions and delineate the scope of patent rights granted to patent holders. These documents are key parts of the patenting process and are publicly accessible typically 18 months after the application or first filing date. The format and content can vary by jurisdiction but normally include the following elements. Fig. 2 displays an example patent document.
Publication information includes the file number and date of patent (application) publication.
@@
Citations are lists of prior arts and other patents referenced in the document or by examiners.
@@
Drawings are visual representations of the invention, disclosing important aspects of the invention as well as embodiments to support the textual description. Fig.3 shows an example patent drawing. Apart from patent texts, researchers also use patent drawings for patent analysis, which is introduced in Appendix A.
@@
Claims define the legal scope of the patent. Each claim is a single sentence, which describes the invention in specific features that make the invention novel and not easily derivable (obvious) from the state of the art (described by any source, not only patent documents).
Figure 3: An example drawing from patent US 10,854,933. Many figures in patents are generic without the corresponding description. Drawings tend to be less generic in patents on pharmaceuticals or mechanics, often consisting of graphs, images, or models. Some drawings are also of poor resolution, pixelation, or low quality. The reference numbers indicate specific elements introduced in the patent description. The reference numbers have to be named by the same term consistently throughout the patent application, which can also substantially deviate from conventions in the field. The reference numbers are typically used for invention features and listed in the claims.
@@ 2.2 Patent Life Cycle
The patent life cycle encompasses several stages, from the initial conception of an invention to its eventual expiration. It can be broadly divided into pre-grant and post-grant phases (Fig. 4).
Figure 4: Patent life-cycle from pre-grant to post-grant stage
Pre-grant Stage. In the beginning, inventors conceptualize, design, and develop their inventions. If inventors hope to obtain patent protection for their inventions, they need to apply to the patent office. To ensure novelty and inventiveness (non-obviousness), inventors may preventively search for existing patents and public disclosures to avoid unnecessary cost and effort in case of prior disclosures. Additionally, patent documents need drafting, describing the invention in detail, including its specifications, claims, abstract, and accompanying drawings. The drafting process typically requires the expertise of a patent professional, such as a patent engineer, attorney, or agent. After the patent application is submitted to the patent office, examiners at the office will screen and evaluate the application for compliance with formal, legal, and technical requirements. This examination typically involves correspondence of the examiner with the inventor or their representation, where clarifications, amendments, or arguments are submitted. [^1] Amendments must not introduce novel features, i.e., inventive content, to avoid losing the original date stamp or the entire file. Notably, an assessment will determine if the invention meets the criteria for patentability, including novelty, inventiveness (non-obviousness), and typically less strictly commercial utility. The requirement of utility may be less strict as offices may see that as the applicant’s problem, except for certain cases. The substantive examination compares the invention with similar documents from the patent literature or any other public source dated earlier, which were found in an initial search by the office. Upon examination and potential resolution of any objections, the patent is either granted—rarely in its original, more frequently in a restricted form based on the identified prior art—or rejected. If the original priority date should be maintained, the invention must not be extended during the examination. [^2] Such additional material should not be added casually but would require proper announcement with a new filing that claims the priority of the base filing. Otherwise, the inadmissibly extended file may be rejected or later challenged in the granted stage.
Post-grant Stage. The granted patent is published, disclosing the details of the invention to the public potentially accompanying a previous application publication. As a remnant of previous pre-computer and pre-AI-search times, the offices classify patents by field of technology for easier search and management. Maintenance fees need to be paid regularly to keep the patent in force. The patent owner can enforce patent rights through legal action when infringement occurs. In addition, third parties can question the validity of patents, and patent owners would need to defend the challenges. Furthermore, companies can analyze patents for technology insights and derive strategies. Finally, the invention enters the public domain after the expiration of patents, allowing anyone to use it without infringement if no other rights cover those aspects.
@@ 2.3 Related Surveys
Patent Analysis. Various surveys in the past summarized certain aspects of patent analysis from a knowledge or procedural perspective. The work of Abbas et al. (2014) represents early research on patent analysis. These early methods included text mining and visualization approaches, which paved the way for future research. Deep learning for knowledge and patent management started more recently. Krestel et al. (2021) identified eight types of sub-tasks attractive for deep learning methods, specifically supporting tasks, patent classification, patent retrieval, patent valuation, technology forecasting, patent text generation, litigation analysis, and computer vision tasks. Some surveys reviewed specific topics of patent analysis. For example, Shalaby and Zadrozny (2019) surveyed patent retrieval, i.e., the search for relevant patent documents, which may appear similar to a web search but has substantially different objectives as well as constraints and has a legally-defined formal character.
@@
Patent Usage. The patent literature records major technological progress and constitutes a knowledge database. It is well known that the translation of methods from one domain to another or the combination of technology from various fields can lead to major innovation. Thus, the contents of the patent literature appear highly attractive for systematic analysis and innovation design. However, the language of modern patents has substantially evolved and diverged from normal technical writing. Recent patent documents are typically hardly digestible for the normal reader and also contain deceptive elements added for increasing the scope or camouflage important aspects. Appropriate data science techniques need to consider such aspects to mine patent databases for engineering design (Jiang et al., 2022b). Data science applied to this body of codified technical knowledge can not only inform design theory and methodology but also form design tools and strategies. Particularly in early stages of innovation, language processing can help with ideation, trend forecasting, or problem–solution matching (Just, 2024).
@@
Patent-Related Language Generation. Whereas the above-listed techniques harvest the existing patent literature for external needs, the strong text-based patent field suggests the use of language models in a generative way. Language processing techniques can, for instance, translate the peculiar patent language to more understandable technical texts, summarize texts, or generate patent texts based on prompts (Casola and Lavelli, 2022). The recent rapid development of generative language models, especially large language models (LLMs) (Zhao et al., 2023) may stimulate more patent-related generation tasks.
@@
Related Topics. A variety of other tasks and applications also have close ties to the patent domain. For example, intellectual property (IP) includes copyrights, trademarks, designs, and a number of more special rights beyond patents. The combined intellectual property literature allows concordant knowledge management, technology management, economic value estimation, and information extraction (Aristodemou and Tietze, 2018). The patent field is closely related to the general legal domain, because they share procedural aspects and the precision of language. Likewise, most patent professionals have substantial legal training or a law degree, and their work often involves legal aspects. Accordingly, NLP techniques in the more general field of law may influence patent-related techniques in the future (Katz et al., 2023).
3 Insights into Patent Texts
Patent text can differ from normal text in multiple aspects, which stimulates a variety of research and entails challenges for the field.
3.1 Long Context
Most research so far has focused on short texts, such as patent abstracts and claims. However, titles and abstracts of patents can be surprisingly generic. Therefore, using patent descriptions that provide comprehensive details and specific embodiment of the invention for patent analysis is highly important, which is neglected by most of the current research. A possible reason is that previous language models cannot handle such long inputs. According to a recently proposed patent dataset (Suzgun et al., 2023), the average number of tokens of a patent description exceeds 11,000, which is longer than the context limit of many previous language models. For example, Llama-2 supports the context length of 4,000 tokens (Touvron et al., 2023). The long context of patent descriptions causes critical challenges for patent analysis.
Notably, researchers have investigated increasing or otherwise handling the context length of LLMs (Chen et al., 2023; Jiang et al., 2023a; Xiong et al., 2023). Xiong et al. (2023), for example, introduced a series of long-context LLMs, which can support the context windows up to 32,768 tokens. Moreover, the state-of-the-art GPT-4 model [^3] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo and Llama-3.1444https://ai.meta.com/blog/meta-llama-3-1/?_fb_noscript=1 even support a context length of 128,000 tokens, which appears promising to process the long patent descriptions. Other LLMs also showed a trend towards long context capabilities, such as Falcon-180B (Almazrouei et al., 2023), Gemini 1.5555https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/ and Claude 3.5666https://www.anthropic.com/news/claude-3-5-sonnet.
3.2 Technical Language
The patent language is highly technical and artificial, including specialized terminology, legal phrases, and sometimes newly coined terms to describe new concepts that may not yet have been widely recognized. Patents regularly define their own terms, which may substantially deviate from everyday language usage and the technical field. Such self-defined terms are often highly artificial and not likely to occur in other relevant documents or even in any dictionary.
Hence, the technical language causes significant challenges for general LLMs for patent analysis, which are trained on normal texts and a large share of more colloquial language. Thus, LLMs may not capture the patent context information effectively, because the important technical terms can be completely new to the LLMs or have different meanings from its pre-training corpora. Embeddings based on distance metrics for synonymity or semantic relationships of terms may not work if terms are defined contrary to their normal use or are entirely new.
3.3 Precision Requirement
The precision requirement and information density of patent texts are higher than in everyday language. [^7] It is worth mentioning that the high importance of legal and formal aspects can tip the balance away from the sciences. Prophetic examples or experiments in patent documents describe anticipated results or theoretical experiments that have not been physically conducted. These are hypothetical scenarios intended to demonstrate the scope of an invention or to anticipate and complicate any competing patent filing for competitors with wild speculations. While many offices only require a basic feasibility of the invention per the claims, such documents can receive incorrect citations and let these prophecies or speculations appear as factual, which can lead to misinformation (Freilich, 2019). The patent language focuses more on precision and accuracy than on readability. Patent texts must be precise and meticulously described to ensure the patent is both defensible and enforceable. Such precision requirement typically leads to high repetitiveness in both terminology and structure of sentences, paragraphs, and sections. Furthermore, sentences are often overburdened because they use relative or adverbial clauses to include specifications for precision or add examples for a wider scope. Additionally, each terminology must be used consistently throughout the document. That means a technical term must not be replaced by other words unless the patent explicitly states that both are identical. In contrast, everyday texts and academic literature tend to vary and paraphrase the wording for better readability. In addition, the patent claim is carefully crafted to define the precise scope and boundaries of the invention’s protection, ensuring that the patent can withstand legal scrutiny.
The precision requirement of patent texts complicates the patent generation tasks because LLMs are likely to generate slightly different words or phrases. Due to the requirement for large quantities of data, most pre-training corpora for LLMs tend to be colloquial and relatively informal. Another smaller portion comes from literary texts, which may be of higher quality but typically prioritize style and linguistic originality over precision and accuracy.
3.4 LLMs for Patent Processing
Patent texts are distinct from everyday texts with respect to long context length, in-depth technical complexity, and high precision to ensure the patent can be granted, defended, and enforced. This type of language often necessitates specific training and experience in patents for accurate reading and interpretation. However, humans usually struggle with names and phrases that conflict with their prior knowledge, or numerous redefined or otherwise complex terms or structures. In contrast, LLMs based on specified training should in principle handle complex structures and newly coined terms that deviate from everyday language, including the specialized vocabulary of experts in the field. Therefore, the recent advancements in LLMs for generative tasks and language processing appear ideally suited for the patent literature.
Despite the apparent fundamental compatibility of LLMs for knowledge extraction and language processing, the application of LLMs in the patent domain remains underdeveloped and not yet highly prominent. Previous studies used word embeddings (e.g., Word2Vec (Mikolov et al., 2013)) and deep learning models (e.g., LSTM (Hochreiter and Schmidhuber, 1997)) for patent analysis tasks. As transformers (Vaswani et al., 2017) showed significant potential in text processing, researchers started to develop transformer-based language models, such as BERT (Devlin et al., 2018) and GPT (Radford et al., 2018). So far, most LLMs adapted and tested in the patent domain are on the level of GPT-2 (Radford et al., 2019) with approximately 1.5 billion parameters. The recent large-sized models with outstanding capabilities have not been extensively investigated in the patent field.
Previous research efforts have included patent analysis, data extraction, and automation of procedures. However, the lack of benchmark tests, such as reference datasets and established metrics, hinders performance evaluation and comparison across different methods. The effectiveness of LLMs depends on the quality of training data. To this end, we have compiled sources and databases of patents with curated datasets tailored for various patent-related tasks in Section 4. Although patent offices have released raw documents for years, publicly available datasets for specific tasks remain scarce. Numerous studies continue to use closed-source data for training and evaluation. Furthermore, patent offices do not provide pre-processed data or broad access to well-structured documents from the process around patents. Although in many countries, patent documents as the disclosure of inventions are considered public domain, they offer only the manual review of individual documents. [^8] This aspect may reflect an important discussion to be led by the field and society. As in other domains, there may be a missing or unclear consensus on which data can be used for machine learning. In stark contrast to other domains, where even books and copyrighted materials may be processed in large numbers, sometimes to create new works that compete with the originals, the patent system is a bit different. The entire modern patent system was developed and is justified as a social contract, trading invention disclosure for temporary exclusion from competition. The descriptions are often public domain and related correspondence with patent offices is made public to ensure transparency. Thus, imposing paywalls would typically violate the mission of patent offices. Admittedly, big-data methods and artificial intelligence were not considerations during the establishment of these legal frameworks. Patent offices may still restrict or complicate larger-scale data access to parts or all contents to avoid commercial exploitation. Some patent offices offer costly paid database access as a stream of revenue, which can be prohibitively expensive for academic research. Important context to know is that in many places patent offices need to fund their own operating expenses (United States Patent and Trademark Office, 2023; President of the European Patent Office and Administrative Council of the European Patent Organisation, 2021; van Pottelsberghe de la Potterie and François, 2009; Frakes and Wasserman, 2014). Therefore, society—which the patent offices are ultimately meant to serve—must reach a consensus on how to manage data access for commercial versus academic purposes, potentially leading to both proprietary and open-source models.
Previous research prominently focuses on short text parts of patents, such as titles and abstracts. However, these texts are typically highly generic and include the least specific texts with little information about the actual invention. These texts also take little time during drafting, making the automated creation of these parts less helpful. By contrast, patent descriptions need to include all details of an invention, and patent claims clearly define the legal boundaries of the invention’s protection, the so-called scope. Since these texts are longer and contain much more useful information, they are worth more attention for both patent analysis and generation.